A 50.0 mL solution of 0.150 M acetic acid (CH3COOH) is titrated with 0.150 M NaOH.

What is the pH after 20.0 mL of base has been added? Ka CH3COOH = 1.8 x 10-5

1
Expert's answer
2022-01-12T21:30:02-0500



"HA+H_2O\rightleftharpoons\>A^-+H_3O^+"


Initial moles of HA (CH"_3" COOH)


"=0.150×\frac{50}{1000}=0.0075\>moles"


Moles of NAOH added

"=0.150×\frac{20}{1000}=0.003\>moles"



Moles of HA remaining "=0.0075-0.003"

"=0.0045\>moles"



Ka"=\frac{[H^+_{aq}][A^-_{aq}]}{[HA_{aq}]}"


"1.8×10^{-5}=\frac{[H^+_{aq}]×0.003}{0.0045}"


"[H^+_{aq}]=2.7\u00d710^{-5}"


PH"=-log\>2.7×10^{-5}"

"=4.5686"